Show that the set P = {a₂t² + a₁t + a₀ |a₂+ a₁ = a₀ and a₂, a₁, a₀ ϵ R} is a group under addition
1
Expert's answer
2020-06-24T19:22:29-0400

Given "P = \{a_2 t² + a_1 t + a_0 | a_2+ a_1 = a_0 \ and \ a_2, a_1, a_0 \in \R\}" .

(i) Closure: Let "A\in P, B \in P" , so "A = a_2 t\u00b2 + a_1 t + a_0 , B = a_2' t\u00b2 + a_1' t + a_0'"

and "a_2+a_1=a_0, a_2'+a_1'=a_0'" .

So, "(a_2+a_2')+(a_1+a_1') = a_0 +a_0'" and "A+B \in P" . Hence, Closure property holds.

(ii) Also, "(A+B)+C=A+(B+C) \ \forall \ A,B,C \in P" .

Hence, Associativity property holds,

(iii) Identity: "\exists \ 0\in P : 0+A = A" for all "A \in P" . Hence, 0 is identity exists in P.

(iv) Inverse: "\forall \ A\in P" , "\exist \ B=-A\in P : A+B=0" .

Hence, the set P is group under addition.